You are trying to overhear a juicy conversation, but from your distance of 20.0 m, it sounds like only an average whisper of 25.0 dB. So you decide to move closer to give the conversation a sound level of 65.0 dB instead. How close should you come? Express your answer in centimeters.
Answer #1

Here,

for the distance

let the distance is d

do = 20 m

as Ifinal = Iinitial - 20 * log(d/do)

65 = 25 - 20 * log(d/20)

solving for d

d = 1/5 m = 0.20 m = 20 cm

the distance shall be 20 cm